OpenStudy (anonymous):

Find the Mistake: Provided below is a solution to the equation 8x + 7 − 2x = 27 − 2. Choose the statement that is true about the given solution. Line # 8x + 7 − 2x = 27 − 2 1.) 6x + 7 = 25 2.) x + 7 = 19 3.) x = 12 A.) The provided solution is wrong. There is a mistake in line #1. B.)The provided solution is wrong. There is a mistake in line #2 C.)The provided solution is wrong. There is a mistake in line #3 D.)The entire solution is correct.

Directrix (directrix):

@turtleluver Is this correct so far? 8x + 7 − 2x = 27 − 2 --------------------- 1.) 6x + 7 = 25